9. A heap of wheat in the form of a cone whose diameter is 10.5 m and height is 3 m. Find iJ.s volume. The heap is to be covered by canvas to protect it from rain. Find the area of the canvas required.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Diameter (d) = 10.6 m Height (h) = 3 m .radius (r) = 10.5 = 6.26 m Volume of t.he heap of wheat = .3!.vh = a1 )( 22 )( 5.26 )( 5.25 )( 3 ms = 86.625 ms. SURFACE AREAS AND VOLUMES 203 Slant height of the cone = (5.25)2 + ca>2 m = 27.5625 + 9 m = .J36.0025 m "" 6.05 m. Area of canvas required = x 5.25 x 6.05 m2 = 99.825 m2•